# A LLM-Powered Automatic Grading Framework with Human-Level Guidelines Optimization

### v1

@@ 5.7. Ablation Studies
Figure 6. Performance comparison between GradeOpt with misconfidence-based and random-based outer batch selection strategies.
@@
(b) Outer Batch Size
@@
Figure 7. Performance of GradeOpt with different outer and inner iteration batch sizes.
Then, we conduct experiments over the sizes of outer batch $|b_{out}|\in\{20,32,64\}$ and inner batch $|b_{in}|\in\{4,6,8\}$, targeting at exploring the influence of those two hyper-parameters on the performance of GradeOpt. From the results in Figure 7(a), we observe an increasing trend in accuracy and kappa as its outer batch size increases. This observation suggests that increasing the number of examples is always beneficial to the final performance. Similarly, from the Figure 7(b), we find an consistent increasing trend of performance as error number increases. Based on these two findings, we can conclude that the larger batch size is likely to bring performance gains to the GradeOpt.
Figure 8. Performance of GradeOpt with different numbers of iterations.
At last, we study how the iterations number impacts accuracy performance. In our experiment, we explore different iteration numbers, ranging from 1 to 5. Whilst increasing iteration on a minibatch, we utilize the early-stopping signal introduced in Section 4.2.2 to carefully monitor overfitting. As shown in Figure 8, increasing iteration with the help of early-stopping signal can effectively lead to higher test accuracy as well as more stable performance. While five iterations produce higher accuracy, due to limited computational resource, we use three iterations as our default setting.
6. Conclusion
@@
This paper explores fully automating guideline optimization to leverage LLM techniques including reflection and prompt engineering to solve ASAG tasks. We innovatively decompose the ASAG procedure into two steps: guideline optimization and grading. Specifically, we set our focus on automatic guideline optimization to avoid the manual efforts of composing a task-optimal guideline. To further prevent labelling a large amount of data, we propose a two-phase ”train and test-adapt” procedure to maximally tune a guideline on a small training set and securely ensure this optimized output is reliable for large-scale grading. The proposed GradeOpt is a multi-agent guideline optimization system that iteratively leads the LLM to reflect on mistakes, learn answer patterns, and make improving modifications. Empirical experiments on a pedagogical dataset have demonstrated the effectiveness of GradeOpt. One future work is to incorporate human intervention as the solution of several limitations we observed during training, such as incorrect error reflection, inferior prompt generation, etc. Meanwhile with the drastic evolution of prompt engineering techniques, our framework can be further enhanced and serve as a valuable evaluation tool for reducing educators’ burden and providing accurate, consistent assessment of learners’ knowledge.

### v2

@@ 5.7 Ablation Studies
Side-by-side bars: misconfidence strategy cohen’s kappa vs. random.
@@
Figure 6: Performance comparison between GradeOpt with misconfidence-based and random-based outer batch selection strategies.
To answer RQ3, we conduct ablation studies to the nested iteration introduced in Section 4.2. We choose to experiment with $Q_{5}$ from $\mathcal{D}_{1}$ as its relatively simple rubric design of $Q_{5}$ makes the shortest guideline prompt, leaving room for GradeOpt to add in its reflective experience as it iteratively learns from $\mathcal{D}_{2}$. Thus, experimenting with $Q_{5}$ can better showcase GradeOpt’s optimization power. With the experimental results shown in the following sections, we demonstrate the effectiveness of each component.
Line plot: GradeOpt accuracy rises as outer-batch size increases,
plateauing after size 8.
@@
Line plot: GradeOpt accuracy rises as inner-batch size increases, plateauing after size 8.
(b) Inner Batch Size
@@
Figure 7: Performance of GradeOpt with different outer and inner iteration batch sizes.
First, we demonstrate the superiority of our misconfidence-based batch sampling strategy by comparing it with the random-based one. From Figure 6, we can observe that misconfidence-based batch sampling results are more consistent and accurate. While random selection generates optimal guidelines in 2 to 4 rounds, misconfidence-based selection consistently optimizes guidelines in 3 rounds. This, together with high predictive accuracy and alignment brought by misconfidence-based selection, makes the system reliable in practical educational scenarios as the required training round number is coherent. Then, we conduct experiments over the sizes of outer batch $|b_{out}|\in\{20,32,64\}$ and inner batch $|b_{in}|\in\{4,6,8\}$, targeting at exploring the influence of those two hyper-parameters on the performance of GradeOpt. From the results in Figure 7, we observe an increasing trend in accuracy and kappa as its outer batch size increases. This observation suggests that increasing the number of examples is always beneficial to the final performance. Similarly, from the Figure 7, we find an consistent increasing trend of performance as error number increases. Based on these two findings, we can conclude that the larger batch size is likely to bring performance gains to the GradeOpt. At last, we study how the iterations number impacts accuracy performance. In our experiment, we explore different iteration numbers, ranging from 1 to 5. Whilst increasing iteration on a minibatch, we utilize the early-stopping signal introduced in Section 4.2.2 to carefully monitor overfitting. As shown in Figure 8, increasing iteration with the help of early-stopping signal can effectively lead to higher test accuracy as well as more stable performance. While five iterations produce higher accuracy, due to limited computational resource, we use three iterations as our default setting.
Influence of iteration number.
Figure 8: Performance of GradeOpt with different numbers of iterations.
6 Case Study
In this section, we conduct case study by going through the refinement history and presenting the extraordinary performance of GradeOpt in optimizing the guidelines to match with task specifications and training dataset. We showcase example responses of Reflector and Refiner separately.
6.1 Reflector
The Reflector reflects and gives feedback when provided with a small selection of error samples. For example, as shown in Table 5, given a selection of errors in the prompt, Reflector follows a two-step reflect-and-suggest approach. It first provides an explanation to each error, illustrating the reason for why the mistake is made and analyzing why the response should be assessed differently. Based on this reflection, Reflector next discusses its suggestion of rule modification, based on the evidence from error sample.
Table 5: Case Study for Reflector
Example Error Teacher Response: “They are equal proportions because they are situations during the same time and the same angle of the sun.” - Expert Label: 2; AI Prediction: 1
Reflector’s Error Feedback Step 1: Explanation of Mistake - The classifier predicted a score of 1 because the response mentions equal proportions and references conditions affecting proportionality (same time and same angle of the sun). However, it lacks specific evidence from the student’s work and a thorough interpretation of the student’s mathematical thinking. The correct label is 2 because the response does provide a clear explanation of the proportional relationship and the conditions under which it holds, which is a nuanced understanding of proportionality.
Step 2: Rule Modification - Rule: If the response explains equal proportions and references conditions affecting proportionality, classify as 2, even if it lacks specific evidence from the student’s work.
6.2 Refiner
As introduced in Section 4.2.1, Refiner is tasked with guideline modification given the feedback from Reflector. Refiner can edit by adding, changing, removing, etc. We first showcase an example of Refiner adding a rule in Table  6, in response to Reflector’s feedback in Section 6.1. We observe that Refiner takes Reflector’s rules suggested in “Step 2: Rule Modification”. The new rule explains two components for grading a certain type of answer: first, the answer pattern, by saying “if the teacher’s response explains equal proportions and references conditions affecting proportionality, …, even if it lacks specific evidence from the student’s work”; second, the score assignment, by saying “classify as 2”. It additionally adds in details from the example by noting “(e.g., same time, same angle of the sun)” and includes the whole answer to give an elucidative grasping of to what answers this rule can be applied. To conclude, when adding rules Refiner is to include new types of answers, concluding their patterns and explaining word or phrase details. Next, we show an example of Refiner editing the guidelines. This happens when Refiner determines that a defective rule can be adjusted to give a better explanation or more complete details. As the example in Table 6 shows, part of the given rule is misleading, so Refiner revises this part. The changed rule first redefines the scenarios when this rule can be utilized, then it provides a detailed explanation by citing the answer and illustrating how the answer falls into the pattern category defined in the new rule to receive the correct score.
Table 6: Case Study for Refiner
Action Refiner’s generated rules
Add “1. Equal Proportions and Conditions:
- Rule: If the teacher’s response explains equal proportions and references conditions affecting proportionality (e.g., same time, same angle of the sun), classify as 2, even if it lacks specific evidence from the student’s work.
- Example: They are equal proportions because they are situations during the same time and the same angle of the sun.”
Edit Before Edit: “- Award 1 point if … explicit evidence. For instance, if the response mentions that the student understands the unit rate, which is related to the concept of proportionality. (misleading statement)
After Edit: “- Award 1 point if … explicit evidence. For example, if the response states that the student might have a limited understanding of proportionality, it should be awarded 1 point. For instance, in the response “their answer makes sense only if there is a proportional relationship between the height of the object and the length of the shadow," teacher mentioned “limited/partial understanding of the proportional relationship" but lacked depth. (ambiguity resolved)
7 Conclusion
@@
This paper explores fully automating guideline optimization to leverage LLM techniques including reflection and prompt engineering to solve ASAG tasks. We innovatively decompose the ASAG procedure into two steps: guideline optimization and grading. Specifically, we set our focus on automatic guideline optimization to avoid the manual efforts of composing a task-optimal guideline. To further prevent labeling a large amount of data, we propose a two-phase “train and test-adapt" procedure to maximally tune a guideline on a small training set and securely ensure this optimized output is reliable for large-scale grading. The proposed GradeOpt is a multi-agent guideline optimization system that iteratively leads the LLM to reflect on mistakes, learn answer patterns, and make improving modifications. Empirical experiments on two pedagogical datasets have demonstrated the effectiveness of GradeOpt.
8 Acknowledgments
This work was supported in part by the National Science Foundation under Grant No. 1813760, No. 2405483, No. 2200757 and No. 2234015. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. We thank Dr. Clare Carlson for help in revising one of the scoring rubrics.